OpenStudy (anonymous):

what is the period of the sinusoid give by y=-2sin(2pi/7x)

OpenStudy (anonymous):

\[-2\sin \frac{ 2\pi }{ 7}x\]

OpenStudy (alekos):

for a standard equation y=asinωx the period T= 2π/ω